# Redistricting Reforms Reduce Gerrymandering by Constraining Partisan Actors

### v2

@@ Abstract
Political actors often manipulate redistricting plans to gain electoral advantages, a process known as gerrymandering. Several states have implemented institutional reforms to address this problem, such as establishing map-drawing commissions. Estimating the impact of such reforms is challenging because each state structures bundles of rules in different ways. We model redistricting as a sequential game, where each state’s equilibrium solution summarizes multi-step institutional interactions as a single-dimensional score. We argue this score measures the leeway political actors have over the partisan lean of the final plan. Using a differences-in-differences design, we demonstrate that reforms reduce partisan bias and increase competitiveness when they constrain partisan actors. We perform a counterfactual policy analysis to estimate the effects of enacting recent reforms nationwide. Though commissions generally reduce bias, reforms that restrict partisan actors in multiple ways like removing veto points (Michigan) are much more effective than commissions where parties retain some control (Ohio).
@@ 1 Introduction
Reform efforts to limit gerrymandering are often designed to constrain partisan map drawers. They include the establishment of independent map-drawing commissions and the introduction of court oversight over proposed plans (Cain,, 2012). Estimating the causal impact of these institutional reforms, however, is challenging for three reasons. The first is the problem of treatment complexity. Redistricting reform efforts must intervene in a complex and multidimensional process. The specific rules governing each state’s redistricting process differ in many ways, including who proposes initial maps and whether or not courts can intervene. Second, states that adopt redistricting reforms may differ in both observable and unobservable ways from states that do not. This selection bias problem is common to any observational study and must be addressed to estimate credible causal effects.
@@
In this paper, we propose a new methodological approach that addresses the above challenges to study the causal impact of redistricting reform. To deal with the treatment complexity, we first collect and standardize information about each state’s relevant laws, starting from the initial map drawer (i.e., the redistricting commission or legislature) and following through varying stalemate processes and opportunities for court intervention. We then reduce the differing districting procedures across states to a single theoretically informed parameter by modeling redistricting reform as a zero-sum sequential game. We use our original dataset of institutional procedures to characterize the players and available moves in the game. Analyzing the game allows us to measure the ability of partisan players to maximize the partisan lean of a redistricting plan. The Nash equilibrium of the game is a measure of the “leeway” that a single party has over the final redistricting plan.
@@
Specifically, we address the selection bias problem by using our leeway measures as a continuous treatment variable in a differences-in-differences (DiD) design applied to the 2010 and 2020 redistricting cycles. For example, states with Democratic control tend to enact more liberal policies (Caughey et al.,, 2017). This approach addresses potential confounding by comparing changes in states that have enacted reforms to those in similar states that have not, under a parallel trends assumption (Callaway et al.,, 2024). We use this model to estimate how changes in the map drawer’s leeway influence resulting plans across a set of both partisan and nonpartisan outcomes, such as the efficiency gap (Stephanopoulos and McGhee,, 2015) and responsiveness to swings in partisan preferences.
Our approach uses a formal model to summarize a multi-dimensional treatment variable, and then apply a causal inference methodology to adjust for bias due to unobserved confounding. However, we do not adopt a structural modeling approach, which would require modeling a state’s decision to adopt particular reforms for their redistricting process. The reason is that, while we are able to measure all the institutional features of redistricting processes, it is impossible to observe all the factors which affect a state’s decision to adopt redistricting reforms. Instead, we use a DiD design to address the issue of unobserved confounding.
@@
We find that commissions can generally reduce existing pro-Republican bias, but the details of commission structure matter. In particular, unlike reforms adopted in Ohio and New York, a Michigan-style nonpartisan commission has no partisan veto points. We find that implementing Michigan-style reforms nationwide leads to an additional 22 Democratic seats, on average. Michigan-style reforms also increase electoral responsiveness whereas those adopted in New York and Ohio do not.
@@
Beyond the substantive findings, this paper advances a new methodology for studying redistricting reform, which we hope will translate to the study of other institutional systems. Reform efforts in redistricting have produced diverse institutional changes across states. While most scholars have classified reforms into different categories (Cain,, 2012; Edwards et al.,, 2017; Warshaw et al.,, 2022; Nelson,, 2023), such an approach may miss important differences in these institutional changes.
@@
In contrast, we use formal modeling to place these complex institutional characteristics on a continuous univariate scale to summarize how they constrain partisan actors. This theoretically driven approach, which we demonstrate accurately predicts empirical patterns, makes it possible to apply a difference-in-differences strategy to our complex setting for credible causal inference.
While our work estimates the causal effects of redistricting reform as a whole, much of existing causal analysis has focused on one part of the process, such as voter registration patterns after the Shelby decision (Komisarchik and White,, 2021) and turnout in packed districts (Fraga et al.,, 2022). Other studies have reported positive associations between the adoption of commissions and competitiveness (Carson and Crespin,, 2004; Carson et al.,, 2014; Best et al.,, 2021; Nelson,, 2023) as well as a negative association between commissions and incumbent protection (Nelson,, 2023). These studies, however, do not adjust for various confounding factors considered in our analysis.
@@ 2 Overview of Redistricting Processes and Reforms
Figure 1: Redistricting procedures for all 44 states with more than one district in 2020. Each vertical column indicates a separate step in the redistricting process, and nodes indicate different procedures that each state can adopt at that step. The width of each area connecting the nodes is proportional to the number of states with that specific combination of procedure at both ends. Yellow nodes indicate actors or institutions that are not explicitly partisan, while blue nodes indicate explicitly partisan actors or choices. Here, we collapse multiple potential stalemate and veto procedures into one step for visual clarity (e.g., Governor + Voters indicates the possibility of a first veto by a governor, and a second by the voters).
As explained in Section 3.1 below, we collect information about the redistricting process used in each state for the 2010 and 2020 cycles. Figure 1 summarizes this dataset and illustrates the diversity of 2020 redistricting procedures across states. For example, 29 states drew initial plans in their legislature, while 15 states used an independent redistricting commission instead. There is significant variation in the procedures following this initial draw, with seven distinct veto mechanisms across all the states. Much of this variation originates from state-specific policy campaigns that have adopted widely different goals driven by local actors (Keena et al.,, 2021). For example, California created a redistricting commission that has the power of drawing congressional district boundaries through a series of ballot propositions in the early 2000s. In recent years, several states have also implemented redistricting reforms in response to organized efforts to limit partisan influence. For example, Michigan voters approved a constitutional amendment to establish an independent redistricting commission in 2018.
@@ 3 A Theoretical Model of Institutional Leeway
To study the impact of redistricting reforms and processes, we first develop a standardized set of 14 institutional features that capture the most important actors in congressional redistricting processes across states. We collect data on these features so that it is possible to compare redistricting processes consistently across states, despite the differences in the laws and bodies governing redistricting in each state.
Next, we develop a sequential redistricting game to summarize these features in a theoretically informed manner. Specifically, we use the Nash equilibrium of this game as a one-dimensional measure of the institutional leeway political actors have over the partisan lean of the resulting redistricting plan. Finally, we empirically validate the proposed measure by demonstrating that it is not particularly sensitive to model specification and predicts redistricting outcomes well. We emphasize that the proposed measure is a theoretically motivated summary of our high-dimensional treatment variable of interest. Since this measure does not use the outcome variable, we are able to apply the standard causal inference methodology to adjust for observed and unobserved confounding factors.
@@ 3.1 Data on the relevant institutional features
Our standardized coding of relevant institutional features is based on the prototypical redistricting process shown in Figure 2. First, an initial map drawer proposes a plan that may be vetoed by other actors. If the plan is not vetoed, it can be challenged in court or not. If it is vetoed, there is another round of map drawing. If a plan is vetoed twice, or if the initial map drawer cannot agree on a plan, then a different institution (often a court) must resolve the stalemate and adopt a plan.
@@ 3.2 The redistricting game
In each state, the nodes in the game tree can be controlled by different parties, or by neither party (e.g., when a supermajority is required to adopt a plan and neither party has supermajority control). Any split-control nodes, as well as a node for state court action, are considered moves by nature. Moves at one node do not affect the set of actions available at other nodes. Some nodes involve discrete choices such as whether to veto a plan or not, others are labeled “choose bias,” meaning that the player at that node draws a plan with a chosen amount of partisan bias favoring either party. This bias is exactly the utility for the party of the chosen plan: a plan with bias $x$ is worth $x$ to the Republicans and $-x$ to the Democrats. We need not quantify exactly what the utility measure is as a function of a specific plan chosen; it suffices to let the parties try to maximize an abstract univariate measure of partisan bias. We let the bias score range from $-4$, indicating a maximum Democratic advantage, to $+4$, indicating a maximum Republican advantage.
For concreteness, consider the case of redistricting in Michigan, as described earlier. The first move belongs to the independent commission. It has to pick the amount of partisan bias $-4<x<4$ in the plan that it adopts. If this plan is ultimately the final enacted plan, Republicans receive utility $x$ and Democrats receive utility $-x$. The stalemate move also belongs to the commission.
If a plan is adopted in the first round, it proceeds to a possible court review. The court may decide to accept a legal challenge, decide in favor of the plaintiffs, and redraw the map; this choice is considered a move by nature. If the court review results in a redrawn map with partisan bias $x^{\prime}$, then the Republicans receive utility $x^{\prime}$ and the Democrats receive utility $-x^{\prime}$. In Michigan, court review is explicitly allowed on partisan grounds and challenges under the federal Voting Rights Act are possible, so there is moderate probability that the commission-adopted plan will be overturned. If no plan is enacted by the commission in the first round, the commission is again tasked with drawing a plan. If the commission fails to enact a plan again, courts must step in and redraw district lines to ensure compliance with federal constitutional “one person, one vote” apportionment requirements. This is also considered a move by nature.
@@
All in all, the game specification depends on 19 parameters which govern the moves by nature, with most of these parameters relating to the court challenge process. Rather than fix these parameters to arbitrary constants, we place a prior distribution on each parameter over a range of probable values. We simulate 100 different draws from this joint prior distribution; each draw generates a slightly different game specification. We then average our results across the random draws. [^6] We average, rather than propagating each draw through the estimation step. This has computational benefits and reduces the complexity of an outcome model (the outcome model would need to account for a different scale of treatment values across different draws). More importantly, however, the variation in the estimated equilibria, as described below, is quite small compared to other sources of uncertainty.
@@ 3.3 Equilibrium solution as a treatment variable
The expected outcome of both the stalemate-then-court and court-alone choices are determined by the parameters governing the moves by nature. The overall equilibrium is a bias of $-0.03$, indicating that there is essentially no bias in favor of either party. This solution process is automated across all of the states and is carried out on each of the 100 different draws from the prior on the game’s parameters.
For our causal analysis, the first treatment measure is exactly the game’s Nash equilibrium, averaged across the prior. This measure depends on which parties control each node in the game tree. We also generate a second treatment measure that does not depend on the current party control. This is calculated by assigning a single party (here, the Democrats [^7] The game is completely symmetric between parties, with one exception: due to racial polarization in the U.S., challenges to redistricting plans under the VRA are much more likely for Republican-leaning plans, and when these challenges prevail the remedy almost always results in a plan more favorable to Democrats. In keeping with the goal of calculating the maximum leeway in each state, we therefore assign states to Democratic control under the counterfactual scenario here. Assigning to Republican control instead would slightly reduce the calculated leeway for those states previously subject to DOJ preclearance.) to control each node in the game tree that belongs to a partisan actor—legislature, governor, or partisan commission—and then recalculating the average Nash equilibrium. We refer to this equilibrium as the maximum leeway of a state’s redistricting process, since it captures the expected bias under a worst-case partisan outcome where all of the levers of state government are controlled by one party. The first treatment measure we refer to as the realized leeway, since it depends on the realized values of the party control for each state institution.

### v3

@@ Abstract
Political actors often manipulate redistricting plans to gain electoral advantages, a process known as gerrymandering. Several states have implemented institutional reforms to address this problem, such as establishing map-drawing commissions. Estimating the impact of such reforms is challenging because each state structures its processes and rules differently. We model redistricting as a sequential game whose equilibrium solution summarizes multi-step institutional interactions as a univariate score. We argue this score measures the leeway political actors have over the partisan lean of the final plan. Using a differences-in-differences design, we demonstrate that reforms reduce partisan bias and increase competitiveness when they constrain partisan actors. We perform a counterfactual policy analysis to estimate the effects of enacting recent reforms nationwide. Though commissions generally reduce bias, reforms that restrict partisan actors in multiple ways, like removing veto points (Michigan), are more effective than commissions where parties retain some control (Ohio).
@@ 1 Introduction
1.1 Methodological challenges and proposed approach
Reform efforts to limit gerrymandering are often designed to constrain partisan map drawers. They include the establishment of independent map-drawing commissions and the introduction of court oversight over proposed plans (Cain,, 2012). Estimating the causal impact of these institutional reforms, however, is challenging for three reasons. The first is the problem of treatment complexity. Redistricting reform efforts must intervene in a complex and multidimensional process. The specific rules governing each state’s redistricting process differ in many ways, including who proposes initial maps and whether or not courts can intervene.
Second, states that adopt redistricting reforms may differ in both observable and unobservable ways from states that do not. These differences include statewide variations in institutional characteristics and political contexts. This confounding bias problem is common to any observational study and must be addressed to estimate credible causal effects.
@@ 1.1 Methodological challenges and proposed approach
Figure 1: Schematic summary of our methodology. Our approach is designed to address three key methodological challenges in the study of redistricting reform. First, we address treatment complexity by modeling the redistricting process as a zero-sum sequential game to estimate theoretically informed parameters that serve as our treatment. Second, we address outcome complexity by generating representative distributions of simulated redistricting plans for each state, which adjust for state-specific changes in political geography. Finally, we address confounding bias in causal effect estimation with a difference-in-differences design that uses simulated alternative to strengthen the credibility of the parallel trends assumption.
In this paper, we propose a new methodological approach that addresses the above challenges to study the causal impact of redistricting reform. Figure 1 summarizes our methodology. To deal with the treatment complexity, we first collect and standardize information about each state’s relevant laws, starting from the initial map drawer (i.e., the redistricting commission or legislature) and following through varying stalemate processes and opportunities for court intervention. We then reduce the differing districting procedures across states to a single theoretically informed parameter by modeling redistricting reform as a zero-sum sequential game. We use our original dataset of institutional procedures to characterize the players and available moves in the game. Analyzing the game allows us to measure the ability of partisan players to maximize the partisan lean of a redistricting plan. The Nash equilibrium of the game is a measure of the “leeway” that a single party has over the final redistricting plan.
@@
Specifically, we address the confounding bias problem by using our leeway measures as a continuous treatment variable in a differences-in-differences (DiD) design applied to the 2010 and 2020 redistricting cycles. For example, states with Democratic control tend to enact more liberal policies (Caughey et al.,, 2017). This approach addresses potential confounding by comparing changes in states that have enacted reforms to those in similar states that have not, under a parallel trends assumption (Callaway et al.,, 2024). We use this model to estimate how changes in the map drawer’s leeway influence resulting plans across a set of both partisan and nonpartisan outcomes, such as the responsiveness to swings in partisan preferences and the number of expected seats per party.
Our approach uses a formal model to summarize a multi-dimensional treatment variable, and then applies a causal inference methodology to adjust for bias due to unobserved confounding. However, we do not adopt a structural modeling approach, which would require modeling a state’s decision to adopt particular reforms for their redistricting process. The reason is that, while we are able to measure all the institutional features of redistricting processes, it is impossible to observe all the factors which affect a state’s decision to adopt redistricting reforms. Instead, we use a DiD design to address the issue of unobserved confounding.
@@ 1.2 Summary of findings
We find that commissions can generally reduce existing pro-Republican bias, but the details of commission structure matter. In particular, unlike reforms adopted in Ohio and New York, a Michigan-style nonpartisan commission has no partisan veto points. We find that implementing Michigan-style reforms nationwide leads to an additional 6.2 Democratic seats, on average. All three styles of reforms increase electoral responsiveness.
1.3 Contributions to the literature
We make two primary substantive contributions to the literature on redistricting reform. First, we help address a debate in the literature about whether, and how, reforms impact redistricting plans. Existing work on the effectiveness of redistricting reforms presents conflicting arguments and findings, with Nelson, (2023) concluding that “the efficacy of redistricting reforms is contested in political science” (p. 207). For example, some studies argue that commissions produce fairer plans, largely by removing self-serving electoral incentives from legislative map drawers (Carson and Crespin,, 2004; Carson et al.,, 2014; Edwards et al.,, 2017; Lindgren and Southwell,, 2013; Litton,, 2012; McDonald,, 2004; Nelson,, 2023; Keena et al.,, 2021). In contrast, others reach much less optimistic conclusions about the efficacy of redistricting commissions and find limited or no impact of reforms on outcomes like competitiveness and partisan bias (Cottrill,, 2012; Henderson et al.,, 2018; Kousser et al.,, 2018; Miller and Grofman,, 2013; Seabrook,, 2017). We contribute to this debate by formalizing and presenting an approach that focuses on how reforms impact leeway—the relative control political actors have over the resulting redistricting plan.
Some work has focused on the long-term decrease of partisan bias in redistricting. Caughey and Warshaw, (2022) highlights how the size of the partisan bias has largely decreased since the 1940s, especially after the reapportionment revolution, though the trend has flattened in recent cycles. We focus on this most recent period. While we do not explain the long-term decrease, we offer policy evaluations on how restrictive reforms could help restart the decrease.
Second, we contribute to the literature on redistricting reform by presenting the most comprehensive empirical evidence yet available about how the full process of redistricting shapes political outcomes. Most existing studies examine reform structures by comparing outcomes from single aspects (e.g., who draws initial maps or whether courts can intervene) of much more complex redistricting process (Carson and Crespin,, 2004; Carson et al.,, 2014; Edwards et al.,, 2017; Nelson,, 2023). We argue that redistricting is best examined by analyzing its entire process. Crucially, our approach allows us to demonstrate that various reforms implemented together can be more effective than a single reform alone.
@@ 1.3 Contributions to the literature
This process-based approach requires several methodological innovations for studying redistricting reform. We believe that the proposed new methodological approach can also be applied to the study of other institutional systems. Reform efforts in redistricting have produced diverse institutional changes across states. While most scholars have classified reforms into different categories (Cain,, 2012; Edwards et al.,, 2017; Warshaw et al.,, 2022; Nelson,, 2023), such an approach may miss important nuances and potential interactions between features of these institutional changes.
@@
In contrast, we use formal modeling to place these complex institutional characteristics on a continuous univariate scale and summarize how they constrain partisan actors. This theoretically driven approach, which we demonstrate accurately predicts empirical patterns, makes it possible to apply a difference-in-differences strategy to our complex setting for credible causal inference.
We also advance the literature on redistricting reform by addressing the aforementioned three methodological challenges in our unified approach. First, most studies of redistricting electoral reforms have ignored treatment complexity by focusing on only one or two aspects of the reform process at a time. For example; some researchers use an indicator for the existence of a redistricting commission (e.g., Carson and Crespin,, 2004; Carson et al.,, 2014); while others account for different types of commissions (e.g., Nelson,, 2023; Edwards et al.,, 2017). Similarly, much of the existing causal research on redistricting has focused on a single aspect of the process, such as being directly impacted by the Shelby decision (Komisarchik and White,, 2021) or being placed in a packed district (Fraga et al.,, 2022). In contrast, we model the entire process of redistricting using a unified formal theoretic framework. This approach enables us to examine how various institutional features affect redistricting outcomes rather than studying each feature in isolation.
Second; much of prior work has been descriptive in nature and primarily relied upon cross-sectional comparisons (e.g., Carson and Crespin,, 2004; Carson et al.,, 2014; Edwards et al.,, 2017; Nelson,, 2023; Keena et al.,, 2021; Best et al.,, 2021; Warshaw et al.,, 2022). This can lead to confounding bias if states that adopted redistricting reforms politically differ from those that did not. Notably, many medium-to-large Democratic states have adopted these reforms. We address confounding bias by examining the over-time changes in institutional rules and employing a differences-in-differences design.
Finally, we address outcome complexity through the use of redistricting simulations that account for factors such as political geography. Specifically, these simulated alternative districts allow us to differentiate the impact of redistricting reforms from that of changes in political geography over time. The period between redistricting cycles in the United States is a full decade, allowing political geography to undergo meaningful changes. Most studies, with some exceptions (e.g., Best et al.,, 2021; Warshaw et al.,, 2022), do not account for changes in underlying political geography when studying redistricting reform. Even studies that address the outcome complexity do not address the above treatment complexity issue.
@@ 2 Overview of Redistricting Processes and Reforms
Figure 2: Redistricting procedures for all $4$ states with more than one district in 2020. Each vertical column indicates a separate step in the redistricting process, and nodes indicate different procedures that each state can adopt at that step. The width of each area connecting the nodes is proportional to the number of states with that specific combination of procedure at both ends. Yellow nodes indicate actors or institutions that are not explicitly partisan, while blue nodes indicate explicitly partisan actors or choices. Green nodes indicate cases where the procedure is not known or does not exist. Here, we collapse multiple potential stalemate and veto procedures into one step for visual clarity (e.g., Governor + Voters indicates the possibility of a first veto by a governor, and a second by the voters).
As explained in Section 3.1 below, we collect information about the redistricting process used in each state for the 2010 and 2020 cycles. Figure 2 summarizes this dataset and illustrates the diversity of 2020 redistricting procedures across states. For example, 29 states drew initial plans in their legislature, while 15 states used an independent redistricting commission instead. There is significant variation in the procedures following this initial draw, with seven distinct veto mechanisms across all the states. Much of this variation originates from state-specific policy campaigns that have adopted widely different goals driven by local actors (Keena et al.,, 2021). For example, California created a redistricting commission that has the power of drawing congressional district boundaries through a series of ballot propositions in the early 2000s. In recent years, several states have also implemented redistricting reforms in response to organized efforts to limit partisan influence. They include Michigan whose voters approved a constitutional amendment to establish an independent redistricting commission in 2018.
@@ 3 A Theoretical Model of Institutional Leeway
To study the impact of redistricting reforms and processes, we must first address the problem of treatment complexity. We develop a standardized set of 14 institutional features that capture the most important actors in congressional redistricting processes across states. We collect data on these features so that it is possible to compare redistricting processes consistently across states, despite the differences in the laws and bodies governing redistricting in each state. This step corresponds to the left top corner of Figure 1.
Using this data, we develop a sequential redistricting game to summarize these features in a theoretically informed manner (see the “treatment” box of Figure 1). Specifically, we use the Nash equilibrium of this game as a one-dimensional measure of the institutional leeway political actors have over the partisan lean of the resulting redistricting plan. Finally, we empirically validate the proposed measure by demonstrating that it is not particularly sensitive to model specification and predicts redistricting outcomes well.
We emphasize that the proposed measure is a theoretically motivated summary of our high-dimensional treatment variable of interest. Since this measure does not use outcome variables, we are able to apply standard causal inference methodology to adjust for observed and unobserved confounding factors.
@@ 3.1 Data on the relevant institutional features
Our standardized coding of relevant institutional features is based on the prototypical redistricting process shown in Figure 3. First, an initial map drawer proposes a plan that may be vetoed by other actors. If the plan is not vetoed, it can be challenged in court. If it is vetoed, there is another round of map drawing. If a plan is vetoed twice, or if the initial map drawer cannot agree on a plan, then a different institution (often a court) must resolve the stalemate and adopt a plan.
@@ 3.2 The redistricting game
In each state, the nodes in the game tree can be controlled by different parties, or by neither party (e.g., when a supermajority is required to adopt a plan and neither party has supermajority control). Any split-control nodes, as well as a node for state court action, are considered moves by nature. Moves at one node do not affect the set of actions available at other nodes. Some nodes involve discrete choices such as whether to veto a plan or not, others are labeled “choose bias,” meaning that the player at that node draws a plan with a chosen amount of partisan bias favoring either party. This bias is exactly the utility for the party of the chosen plan: a plan with bias $x$ is worth $x$ to the Republicans and x-x to the Democrats. We need not quantify exactly what the utility measure is as a function of a specific plan chosen; it suffices to let the parties try to maximize an abstract univariate measure of partisan bias. We let the bias score range from 4-4, indicating a maximum Democratic advantage, to +4+4, indicating a maximum Republican advantage.
For concreteness, consider the case of redistricting in Oregon. In 2020, the first move belonged to the Democratic controlled legislature. It has to pick the amount of partisan bias 4<x<4-4<x<4 in the plan that it adopts. If this plan is ultimately the final enacted plan, Republicans receive utility $x$ and Democrats receive utility x-x. If the legislature fails to adopt a plan, the first stalemate move in Oregon belongs to a nonpartisan commission.
If a plan is adopted in the first round in Oregon, it proceeds to face a potential veto by the Democratic governor or a possible court review. The court may decide to accept a legal challenge, decide in favor of the plaintiffs, and redraw the map; this choice is considered a move by nature. If the court review results in a redrawn map with partisan bias $x^{\prime}$, then the Republicans receive utility $x^{\prime}$ and the Democrats receive utility $x^{\prime}$. In Oregon, court review is explicitly allowed on partisan grounds and challenges under the federal Voting Rights Act are possible, so there is moderate probability that the commission-adopted plan will be overturned. If no plan is enacted by the commission in the first round, the commission is again tasked with drawing a plan. If the commission fails to enact a plan again, courts must step in and redraw district lines to ensure compliance with federal constitutional “one person, one vote” apportionment requirements. This is also considered a move by nature.
@@
All in all, the game specification depends on 19 parameters which govern the moves by nature, with most of these parameters relating to the court challenge process. Rather than fix these parameters to arbitrary constants, we place a prior distribution on each parameter over a range of probable values. We simulate 200 different draws from this joint prior distribution; each draw generates a slightly different game specification. We then average our results across the random draws. [^6] We average, rather than propagating each draw through the estimation step. This has computational benefits and reduces the complexity of an outcome model (the outcome model would need to account for a different scale of treatment values across different draws). More importantly, however, the variation in the estimated equilibria, as described below, is quite small compared to other sources of uncertainty.
@@ 3.3 Equilibrium solution as a treatment variable
For our causal analysis, we use three treatment measures. The first treatment measure is exactly the game’s Nash equilibrium, averaged across the prior. This measure depends on which parties control each node in the game tree. In our analysis of nonpartisan outcomes, we also use the absolute value of the equilibrium to capture the magnitude but not the sign of the expected bias. Finally, we also generate a third treatment measure that does not depend on the current party control. This is calculated by assigning a single party (here, the Democrats [^7] The game is completely symmetric between parties, with one exception: due to racial polarization in the U.S., challenges to redistricting plans under the VRA are much more likely for Republican-leaning plans, and when these challenges prevail the remedy almost always results in a plan more favorable to Democrats. In keeping with the goal of calculating the maximum leeway in each state, we therefore assign states to Democratic control under the counterfactual scenario here. Assigning to Republican control instead would slightly reduce the calculated leeway for those states previously subject to DOJ preclearance.) to control each node in the game tree that belongs to a partisan actor—legislature, governor, or partisan commission—and then recalculating the average Nash equilibrium. We refer to this equilibrium as the maximum leeway of a state’s redistricting process, since it captures the expected bias under a worst-case partisan outcome where all of the levers of state government are controlled by one party. The first and second treatment measures we refer to as the realized leeway and absolute realized leeway, since they depend on the realized values of the party control for each state institution.
3.4 Realized and maximum leeway scores